# Scalable Graph Condensation with Evolving Capabilities

### v1

@@ 5.2.2. Efficiency Comparison
Comparing GECC with the baselines in Figure 2, it is evident that GECC exhibits superior efficiency and scalability. It maintains stable performance while effectively managing computational resources as the graph evolves. Although certain model-based GC methods such as GEOM may slightly outperform GECC at specific time steps, GECC achieves over 100 times faster condensation time and demonstrates a significantly slower increase in computational overhead. For results in more datasets, please refer to Appendix C.
@@ 5.3.1. Impact of Feature Propagation
As feature propagation (Equation 4.3.1) is crucial for generating informative node representations, we compare GECC against an ablated version without feature propagation (labeled w/o propagation) across all time steps. Specifically, for w/o propagation, we set $\alpha_{0}=1$ and all other $\alpha$ coefficients to 00, keeping all other hyperparameters identical. Each experiment is repeated ten times, and we report the average results.
@@ 5.3.2. Impact of Incremental $k$-Means++
Figure 4. Comparison between GECC and GECC without incremental $k$-means++. The boxplots show the distribution (mean and quartiles) of the number of iterations required for clustering.
@@ 5.3.3. Relation between GC and clustering objective
The balanced Sum of Squared Errors (SSE) is a key contribution derived from our theoretical analysis. To assess its impact, we perform an ablation study comparing the performance of GECC with and without repetitive clustering. As shown in Table 4, for instance, on the Citeseer dataset, applying repetitive clustering to select the lowest SSE leads to an absolute performance improvement of 2.7%. These results clearly show that a lower SSE consistently correlates with higher test accuracy.
@@ 6. Conclusion and Outlook
In this study, we address the challenge of evolving graph condensation. We observe that a universal clustering framework can naturally optimize the assignment matrix, thereby achieving the common objectives of existing GC methods. Additionally, we propose a novel balanced SSE metric that further tightens the upper bound of these objectives. In the evolving setting, we find that our clustering approach can be readily adapted to an incremental version, termed incremental $k$-means++. Experimental results demonstrate that balanced SSE improves the performance of clustering-based GC, and incremental $k$-means++ significantly reduces the number of iterations, thereby enhancing efficiency in evolving environments. Future work includes developing more efficient and scalable clustering techniques, especially soft clustering algorithms for larger graph datasets and adaptively optimizing multi-hop weights, which could be beneficial when the graph keeps evolving over time.
@@ Appendix B Method Pipeline
Appendix C Experimental Details
@@
Ogbn-arxiv 169,343 1,166,243 90,941 29,799 48,603 40 Transductive
@@
C.1. Dataset Statistics
In line with most GC studies, we utilize seven datasets in total: five transductive datasets—Citeseer, Cora (Kipf and Welling, 2016), Pubmed (Namata et al., 2012), Ogbn-arxiv, and Ogbn-products (Hu et al., 2020)—and two inductive datasets, Flickr and Reddit (Zeng et al., 2019). Each graph is randomly split, ensuring a consistent class distribution. The details of the datasets statistics are shown in Table 5. We list all evolving information in Table 6, rows above the midline correspond to smaller datasets, and rows below it correspond to larger ones. Reduction rate $r$ is defined as (#nodes in synthetic set)/(#nodes in training set) while $r_{w}$ is (#nodes in synthetic set)/(#nodes of whole graph visible in training stage). The whole graph visible in training stage means the full graph dataset for transductive setting but only the training graph for inductive setting.
@@ C.1. Dataset Statistics
C.2. Platform and Hardware Information
To efficiently execute the clustering algorithm, we run it on Intel(R) Xeon(R) Platinum 8260 CPUs @ 2.40GHz using NumPy (Harris et al., 2020), while the downstream GNN evaluations are conducted on a cluster equipped with a mix of Tesla A100 40GB/V100 32GB GPUs for large datasets and K80 12GB GPUs for small datasets. All GNN models are implemented using the PyG package (Fey and Lenssen, 2019).
C.3. Baselines Selection
@@
C.4. Implementation Details for Variants of GCond
As mentioned in Section 1 and illustrated in Table 1, adapting GCond to an evolving setting is challenging. We employ the structure-free variant of GCond, i.e., GCondX, for easier adaptation, as designing a specific growth mechanism for the condensed graph is nontrivial and requires significant effort. In addition, to manifests the convergence speed difference between GCond and GCond-Init, we implement an early stopping criterion with a patience of 3 during intermediate evaluations. If no improvement in validation performance is observed over 3 consecutive evaluations, the condensation process is terminated.
C.5. Hyperparameters
Compared to existing work and benchmarks in GC, we perform a moderate hyperparameter search on validation set, as detailed in Section 2. The final results are presented in Table 9. During hyperparameter optimization (HPO), we observe that inheriting clustering centroids results in an approximate 1% absolute performance drop for Flickr and Ogbn-arxiv. Therefore, we also treat the use of incremental $k$-means++ as a tunable hyperparameter. Additionally, some datasets do not perform well with a single hyperparameter configuration. To address this, we employ two distinct hyperparameter sets tuned on the first and last time steps, respectively, and select the better-performing one during the evolution process. These two sets are represented using a ”/” separator and are indicated as ”if Dual” when this technique is applied. For all baselines, we use the best hyperparameters reported in their respective papers, as implemented in GC4NC (Gong et al., 2024).
The optimal hyperparameters also offer meaningful insights. First, during the early evolution stage, graphs exhibit higher heterophily compared to later stages. For example, on the Cora dataset, $\alpha_{1}=-0.3$ in the early phase contrasts with $\alpha_{1}=0.9$ later. This pattern likely arises because, in the early stages of a graph, groups have not yet formed; links appear more randomly, making it challenging for nodes to link to similar counterparts. Second, it is noteworthy that some datasets do not rely on second-hop information. This observation is contrary to previous studies (Wu et al., 2019; Luo et al., 2024) that recommend using at least 2-hop propagation. We conjecture that the representation clustering process itself acts as an additional step of feature propagation. Finally, weight decay emerges as a critical factor for the performance of downstream models, suggesting that future work should pay closer attention to its optimization.
@@ C.5. Hyperparameters
Table 9. Optimal hyperparameter settings for each dataset. The columns indicate whether incremental k-means++ (”If Incre.”) and dual HPO (”If Dual”) were used (True/False), along with the best values for the fuzziness parameter, number of repetitions, the $\alpha$ coefficients, and weight decay (”wd”).
Dataset If Incre. If Dual Fuzziness # Rep. $\alpha_{1}$ $\alpha_{2}$ $\alpha_{0}$ wd
@@
Ogbn-arxiv False False 1 1 0.5 0 0.6 1e-3
@@
Appendix D Additional Results
D.1. Performance and Efficiency

### v2

@@ 5.2.2. Efficiency Comparison
Comparing GECC with the baselines in Figure 2, it is evident that GECC exhibits superior efficiency and scalability. It maintains stable performance while effectively managing computational resources as the graph evolves. Numerical experiments show sublinear runtime growth: a fitted power-law exponent (0.3$\approx$ 0.3≈ 0.3) indicates reasonably good scalability. Although certain model-based GC methods such as GEOM may slightly outperform GECC at specific time steps, GECC achieves over 100 times faster condensation time and demonstrates a significantly slower increase in computational overhead. For results in more datasets, please refer to Appendix E.
@@ 5.3.1. Impact of Feature Propagation
As feature propagation (Equation 4.3.1) is crucial for generating informative node representations, we compare GECC against an ablated version without feature propagation (labeled w/o propagation) across all time steps. Specifically, for w/o propagation, we set $\alpha_{0}=1$ and all other $\alpha$ coefficients to 0, keeping all other hyperparameters identical. Each experiment is repeated ten times, and we report the average results.
@@ 5.3.2. Impact of Incremental $k$-Means++
Figure 4. Ablation study. Boxplots show the mean and quartiles of the number of iterations required for clustering.
@@ 5.3.3. Relation between GC and clustering objective
The balanced Sum of Squared Errors (SSE) is a key contribution derived from our theoretical analysis. To assess its impact, we perform an ablation study comparing the performance of GECC with and without repetitive clustering. As shown in Table 4, for instance, on the Citeseer dataset, applying repetitive clustering to select the lowest SSE leads to an absolute performance improvement of 2.7%. These results show that a lower SSE correlates with higher test accuracy.
@@ 6. Conclusion and Outlook
In this study, we address the challenge of evolving graph condensation. We observe that a universal clustering framework can naturally optimize the assignment matrix, thereby achieving the common objectives of existing GC methods. Additionally, we propose a novel balanced SSE metric that further tightens the upper bound of these objectives. In the evolving setting, we find that our clustering approach can be readily adapted to an incremental version, termed incremental $k$-means++.
Experimental results demonstrate that balanced SSE improves the performance of clustering-based GC, and incremental $k$-means++ significantly reduces the number of iterations, thereby enhancing efficiency in evolving environments. Future work includes developing more efficient and scalable clustering techniques, especially soft clustering algorithms for larger graph datasets and adaptively optimizing multi-hop weights, which could be beneficial when the graph keeps evolving over time.
@@ Appendix B Method Pipeline
Appendix C Symbol Table
Table 5. Notations
Symbol Description
General Graph Notations
$G,G_{t}$ An static graph and a snapshot of an evolving graph at time step $t$.
$G^{\prime},G^{\prime}_{t}$ The condensed graph from the static graph and the evolving graph at time step $t$.
$\mathcal{V},\mathcal{E}$ The set of nodes and edges in a graph.
$N,E$ The total number of nodes and edges.
$\mathbf{X},\mathbf{X}_{t}$ The node feature matrix of the static graph and evolving graph at time $t$.
$\mathbf{X}^{\prime},\mathbf{X}^{\prime}_{t}$ The feature matrix of the condensed graph (i.e., the centroids).
$\mathbf{A},\mathbf{A}_{t}$ The adjacency matrix of the original graph and at time $t$.
$\mathbf{A}^{\prime},\mathbf{A}^{\prime}_{t}$ The adjacency matrix of the condensed graph (set to $\mathbf{I}$).
$\mathbf{y},\mathbf{Y},\mathbf{Y}_{t}$ The numeric and one-hot encoded label vectors/matrices of the original graphs.
$\mathbf{Y}^{\prime},\mathbf{Y}_{t}^{\prime}$ The one-hot encoded label of the condensed nodes.
$c$ The total number of classes.
$\mathbf{L},\mathbf{D}$ The graph Laplacian and degree matrices.
$\mathcal{F}$ The update function to generate a new condensed graph.
Evolving Graph Condensation
$B_{t}$ A batch of new graph data arriving at time step $t$.
$m$ The total number of time steps in the evolving sequence.
$r$ The reduction rate, determining the size of the condensed graph.
GECC Algorithm Notations
$\tilde{A}_{t},\hat{A}_{t}$ The adjacency matrix with self-loops and its normalized version.
$K$ The maximum depth for feature propagation.
$F_{t}$ The propagated node feature matrix at time $t$.
$\alpha_{k}$ Learnable weight for combining features from propagation depth $k$.
$n_{t,k}$ The number of nodes belonging to class $k$ at time $t$.
$M,M_{k}$ The total number of condensed nodes (clusters) and per class.
$P_{t}$ The cluster assignment matrix at time $t$.
$C_{t},C_{t-1}$ The set of cluster centroids at time $t$ and $t-1$.
$J(\cdot,\cdot)$ The balanced Sum of Squared Errors (SSE) objective function.
$D(x_{i},C)$ The minimum Euclidean distance from a node $x_{i}$ to a set of centroids $C$.
Algorithm 1 GECC Node Condensation Procedure
1:Adjacency matrix $A_{t}$, feature matrix $X_{t}$, label vector $y_{t}$, reduction rate $r$, label matrix $Y^{\prime}_{t}\leftarrow$ predefined using class distribution in $Y_{t}$
2:Condensed node representations $X^{\prime}_{t}$
3:// Feature Propagation (Eq. 5)
[…]
7:for $k=0$ to $K$ do
8:  $F_{t,k}\leftarrow\hat{A}_{t}^{k}X_{t}$
9:end for
[…]
11:// Class-wise Clustering Setup
12:for each class $k=1,\ldots,c$ do
13:  $n_{t,k}\leftarrow$ number of labeled nodes in class $k$
14:  $M_{k}\leftarrow\lfloor n_{t,k}\times r\rfloor$triangleright Eq. before clustering
15:end for
[…]
17:// Cluster Assignment Matrix $P_{t}$ and Balanced SSE (Eq. 5)
18:for each clustering repetition do
19:  Apply $k$-means or fuzzy $c$-means on $F_{t}$ to get $P_{t}\in\mathbb{R}^{n_{t}\times M}$
20:  Compute cluster centroids: $C_{t}=D_{P_{t}}^{-1}P_{t}^{\top}F_{t}$triangleright Eq. 6
21:  Compute balanced SSE loss:
$J(P_{t},C_{t})=\|F_{t}-P_{t}C_{t}\|^{2}+\|P_{t}^{\top}\mathbf{1}-u\|^{2}$
22:end for
23:Choose clustering with lowest $J(P_{t},C_{t})$
24:// Probabilistic Centroid Initialization (Diverse Coverage)
25:Initialize ${}C_{t-1}=\{\}$
26:for each class $c$ and node $x_{i}\in F_{t}$ do
27:  $D(x_{i},C_{t-1})=\min_{c\in C_{t-1}}\|x_{i}-c\|_{2}$
28:end for
29:Sample new centroid $x_{i}$ with probability:
$P(x_{i})=\frac{D(x_{i},C_{t-1})^{2}}{\sum_{j}D(x_{j},C_{t-1})^{2}}$
[…]
31:// Final Assignment
32:for each node $x_{i}$ do
33:  $\text{Cluster}(x_{i})=\arg\min_{c\in C_{t}}\|x_{i}-c\|_{2}$
34:end for
[…]
Appendix D Complexity analysis
Appendix E Experimental Details
@@
Ogbn-arxiv (-real) 169,343 1,166,243 90,941 29,799 48,603 40 Transductive
@@
Figure 6. Class distribution across five splits of the Ogbn-arxiv-real dataset.
E.1. Dataset Statistics
In line with most GC studies, we utilize seven datasets in total: five transductive datasets—Citeseer, Cora (Kipf and Welling, 2016), Pubmed (Namata et al., 2012), Ogbn-arxiv, and Ogbn-products (Hu et al., 2020)—and two inductive datasets, Flickr and Reddit (Zeng et al., 2019). Each graph is randomly split, ensuring a consistent class distribution. The details of the datasets statistics are shown in Table 6. We list all evolving information in Table 8, rows above the midline correspond to smaller datasets, and rows below it correspond to larger ones. The real-world evolving information of Ogbn-arxiv-real is shown in Table 7. The distribution of the number of nodes in each class is shown in Figure 6.
Table 7. Training Set Splits of Ogbn-arxiv-real by Year Range and Node Count
Split No. Year Range Number of Nodes
1 1971 – 2007 4,980
2 2008 – 2010 7,994
3 2011 – 2012 10,862
4 2013 – 2013 8,135
5 2014 – 2017 58,970
Reduction rate $r$ is defined as (#nodes in synthetic set)/(#nodes in training set) while $r_{w}$ is (#nodes in synthetic set)/(#nodes of whole graph visible in training stage). The whole graph visible in training stage means the full graph dataset for transductive setting but only the training graph for inductive setting.
@@ E.1. Dataset Statistics
E.2. Platform and Hardware Information
To efficiently execute the clustering algorithm, we run it on Intel(R) Xeon(R) Platinum 8260 CPUs @ 2.40GHz using NumPy (Harris et al., 2020), while the downstream GNN evaluations are conducted on a cluster equipped with a mix of Tesla A100 40GB/V100 32GB GPUs for large datasets and K80 12GB GPUs for smaller datasets. All GNN models are implemented using the PyG package (Fey and Lenssen, 2019).
E.3. Baselines Selection
@@
E.4. Implementation Details for Variants of GCondX
As mentioned in Section 1 and illustrated in Table 1, adapting GCondX to an evolving setting is challenging. The experimental results indicate that, although there is a slight efficiency improvement, there is a significant performance drop due to inheriting previous condensation results. To highlight the difference in speed, we additionally implement an early stopping criterion with a patience of 3 for intermediate evaluation; that is, if no improvement in validation performance is observed for 3 consecutive evaluations, the condensation process is terminated.
E.5. Hyperparameters
Compared to existing work and benchmarks in GC, we perform a moderate hyperparameter search on validation set, as detailed in Section 1. The final results are presented in Table 11. During hyperparameter optimization (HPO), we observe that inheriting clustering centroids results in an approximate 1% absolute performance drop for Flickr and Ogbn-arxiv. Therefore, we also treat the use of incremental $k$-means++ as a tunable hyperparameter. Additionally, certain datasets do not perform well under a single hyperparameter configuration. To address this, we employ two distinct hyperparameter sets tuned on the first and last time step, and select the better-performing one during the evolving. The two sets of hyperparameters are represented split by ”/”. We do not tune the hyperparameters for Ogbn-arxiv-real and simply copy it from Ogbn-arxiv. For all baselines, we use the best hyperparameters reported in their respective papers, as implemented in GC4NC (Gong et al., 2024).
Table 11 reveals that the optimal hyperparameters offer meaningful insights. First, during the early evolution stage, graphs exhibit higher heterophily compared to later stages. For example, on the Cora dataset, $\alpha_{1}=-0.3$ in the early phase contrasts with $\alpha_{1}=0.9$ later. This pattern likely arises because, in the early stages of a graph, groups have not yet formed; links appear more randomly, making it challenging for nodes to find similar counterparts. Second, it is noteworthy that some datasets do not rely on second-hop information. This observation is contrary to previous studies (Wu et al., 2019; Luo et al., 2024) that recommend using at least 2-hop propagation. We conjecture that the representation clustering process itself acts as an additional step of feature propagation. Finally, weight decay emerges as a critical factor for the performance of downstream models, suggesting that future work should pay closer attention to its optimization.
@@ E.5. Hyperparameters
Table 11. Optimal hyperparameter settings for each dataset. The columns indicate whether incremental k-means++ (”If Incre.”) and dual HPO (”If Dual HPO”) were used (True/False), along with the best values for the fuzziness parameter, number of repetitions, the $\alpha$ coefficients, and weight decay (”wd”).
Dataset If Incre. If Dual HPO Fuzziness # Rep. $\alpha_{1}$ $\alpha_{2}$ $\alpha_{0}$ wd
@@
Ogbn-arxiv (-real) False False 1 1 0.5 0 0.6 1e-3
@@
Appendix F Additional Results
F.1. Performance and Efficiency
@@
F.2. Why We Choose the Depth=2 ?
We pick $K=2$ based on findings from most deeper GNN papers. Below is a snippet for Cora, Ogbn-arxiv, and Reddit, showing that $K=2$ performs best. We add the following sensitivity analysis.
Table 12. Sensitivity Analysis of GECC Performance by Feature Propagation Depth. The best-performing depth for each dataset is highlighted in bold.
Dataset Feature Propagation Depth
0 1 2 3 4 5
Cora 19.24 78.13 82.99 82.15 81.80 81.41
Ogbn-arxiv 63.64 66.60 66.71 66.13 66.24 66.29
Reddit 90.44 90.44 91.37 91.35 91.28 91.14
F.3. Visualization of Traceability
Our t-SNE-based experiment color-codes original nodes by class and links them to their centroid(s), revealing:
  • Clear class-wise alignment: Synthetic nodes (cluster centroids) align well with the clusters of original nodes, indicating that GECC preserves the semantic structure of each class during condensation.
  • Compact cluster coverage: Each synthetic node is well-embedded within the local neighborhood of its corresponding class, highlighting that the condensed representations accurately cover the original data distribution.
  • Few synthetic outliers: Unlike many condensation approaches that generate scattered or poorly positioned centroids, our method yields minimal outlier centroids, suggesting robust and stable clustering even in high-dimensional spaces.
  • Scalability and consistency: Even in large-scale datasets, our synthetic nodes remain representative and well-distributed, showing the scalability of our clustering method without sacrificing membership fidelity.
Figure 8. T-SNE results of the GECC synthetic node embeddings and original node embeddings across different classes.
Appendix G Discussions
G.1. How the Theory Extend to Non-linear GNNs?
While our theoretical analysis focuses on SGC, we believe local approximations could generalize our approach to non-linear models. However, non-linearities introduce complexities like non-convex optimization and intricate layer interactions, making closed-form guarantees difficult. Many prior works (Gao et al., 2024c; Jin et al., 2022) similarly rely on linear or quasi-linear assumptions. We aim to broaden our framework to fully accommodate non-linear GNNs in future work.
G.2. Clarification of Our Evolving Setting Compared to Incremental Learning
Both our work and incremental learning work such as OpenGC (Gao et al., 2024b) address evolving datasets, yet the settings differ in important ways. Specifically, our paper focuses on batch-wise evolution, where the training set is expanded or updated in batches from the same data source. By contrast, OpenGC deals with a task-wise evolution scenario and is centered on out-of-distribution issues.
G.3. When to Use Hard or Soft Clustering in GECC?
Hard clustering is faster (each node has a single cluster) and suits large datasets or real-time scenarios. Soft clustering allows fractional memberships but is more costly. As more efficient soft clustering emerges, it may become viable for large graphs and potentially improve GECC.